OpenStudy (anonymous):

how to identify the axis of symmetry, the coordnates of the vertex of y=x to the scond power-4

OpenStudy (mertsj):

y=x^2-4 is a parabola. Its vertex is (0,-4) so its aos is x = 0

OpenStudy (anonymous):

so is the axis ofsymmtry always gonnabe 0?

OpenStudy (mertsj):

No. The axis of symmetry always goes through the vertex. If the vertex would be (3,5) the the aos would be x = 3

OpenStudy (anonymous):

how do you find the veretex

OpenStudy (mertsj):

make the equation look like this: y =ax^2 + bx +c yours would be y = 1x^2 +0x -4 Then the x coordinate of the vertex is the opposite of b over 2a. In your case it would be the opposite of 0 over 2, which, of course, is 0. Then find y by plugging 0 into the equation in place of x. The result will be -4. That is the y coordinate of the vertex.

OpenStudy (cathyangs):

the vertex is -b/2a, for y=ax^2+bx+c

OpenStudy (cathyangs):

meant the axis of symmetry XD

OpenStudy (anonymous):

so if the question was y=-x^2+4x-1 the axis of symmetry would be -4?

OpenStudy (anonymous):

or -2